An herbalist has 40 oz of herbs costing $3 per ounce. How many ounces of herbs costing $2 per ounce should be mixed with these 40 oz of herbs to produce a mixture costing $2.50 per ounce?
Answer by stanbon(75887)

An herbalist has 40 oz of herbs costing $3 per ounce. How many ounces of herbs costing $2 per ounce should be mixed with these 40 oz of herbs to produce a mixture costing $2.50 per ounce?
-----------------------------------
Equation:
value + value = value
3*40 + 2x = 2.5*40
120 + 2x = 100
2x = 20
x = 10 oz (# of ounces of $2 herbs needed)
==============================================
Cheers,
Stan H.